OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method. (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 4z = –9 2x + y + 3z = 6 x – 4y + 3z = –9

OpenStudy (texaschic101):

I would use the elimination method because there is 3 unknown variables. That is just my preference. You could use substitution or matrices also. x - 3y + 4z = -9 -->(-2) 2x + y + 3z = 6 -------------- -2x + 6y - 8z = 18 (result of multiplying by -2) 2x + y + 3z = 6 --------------add 7y - 5z = 24 2x + y + 3z = 6 x - 4y + 3z = -9 -->(-2) ----------------- 2x + y + 3z = 6 -2x + 8y - 6z = 18 (result of multiplying by -2) -----------------add 9y - 3z = 24 7y - 5z = 24 -->(3) 9y - 3z = 24 -->(-5) ---------------- 21y - 15z = 72 (result of multiplying by 3) -45y + 15z = - 120 (result of multiplying by -5) ----------------add - 24y = - 48 y = -48/-24 y = 2 7y - 5z = 24 7(2) - 5z = 24 14 - 5z = 24 -5z = 24 - 14 -5z = 10 z = -2 x - 3y + 4z = -9 x - 3(2) + 4(-2) = -9 x - 6 - 8 = -9 x - 14 = -9 x = -9 + 14 x = 5 check... 2x + y + 3z = 6 2(5) + 2 + 3(-2) = 6 10 + 2 - 6 = 6 12 - 6 = 6 6 = 6 (correct) x = 5, y = 2, and z = -2